# QWENDY: Gene Regulatory Network Inference Enhanced by Large Language Model and Transformer

### v2

@@ Abstract
Knowing gene regulatory networks (GRNs) is important for understanding various biological mechanisms. In this paper, we present a method, QWENDY, that uses single-cell gene expression data measured at four time points to infer GRNs. Based on a linear gene expression model, it solves the transformation for the covariance matrices. Unlike its predecessor WENDY, QWENDY avoids solving a non-convex optimization problem and produces a unique solution. Then we enhance QWENDY by the transformer neural networks and large language models to obtain two variants: TEQWENDY and LEQWENDY. We test the performance of these methods on two experimental data sets and two synthetic data sets. TEQWENDY has the best overall performance, while QWENDY ranks the first on experimental data sets.
@@ 1 Introduction
One gene can activate or inhibit the expression of another gene. Genes and such regulation relations form a gene regulatory network (GRN). For $n$ genes, the corresponding GRN is commonly expressed as an $n\times n$ matrix $A$, where $A_{ij}>0/=0/<0$ means gene $i$ has positive/no/negative regulation effect on gene $j$, and its absolute value represents the regulation strength. If the regulation relations of concerned genes are known, one can understand the corresponding biological process or even control it with gene perturbation. Therefore, knowledge of GRNs can be useful in developmental biology [1, 2, 3], and even in the study of macroscopic behavior [4, 5, 6]. Since it is difficult to determine the GRN directly, the common practice is to infer the GRN from gene expression data.
@@
For such single-cell gene expression data measured at multiple time points, although it is the most informative data type under current technology, there are only a few GRN methods developed specifically for this data type [8]. We have developed WENDY method to infer the GRN with this data type [9]. It needs measurement at two time points, and solves the transformation between the covariance matrices of two probability distributions, where the transformation is determined by the GRN. However, it needs to solve a non-convex optimization problem, which has infinitely many solutions, and WENDY will output one solution, determined by the numerical solver.
In this paper, we present an improved version of WENDY that needs data measured at four time points, and it is also based on solving the transformation between covariance matrices. This new method is named QWENDY, where Q stands for quadruple. With the help of data from more time points, QWENDY can uniquely determine the GRN. Besides, QWENDY does not need to conduct non-convex optimizations, but just matrix decompositions.
WENDY method is derived from a simplified gene expression model, and the performance of WENDY is not satisfactory when the model does not fit with experiments. Instead of developing a more complicated gene expression model, we developed TRENDY method [10], which trains a transformer model, a deep learning architecture that can be applied to various fields [11], to transform the input data to better fit the gene expression model. Then the transformed data will produce more accurate GRNs by WENDY. Such more accurate GRNs will be further enhanced by another transformer model.
We test the performance of QWENDY, TEQWENDY, and LEQWENDY on two experimental data sets and two synthetic data sets. The TRENDY paper [10] has tested 16 methods on the same data sets, meaning that we can compare them with our three new methods. In all 19 GRN inference methods, TEQWENDY has the best overall performance. If we restrict ourselves to two experimental data sets, QWENDY ranks first. Therefore, we recommend applying QWENDY or TEQWENDY.
Section 2 reviews other GRN inference methods, especially those related to LLMs. Section 3 introduces QWENDY method. We develop TEQWENDY and LEQWENDY methods in Section 4. Section 5 tests the performance of different methods on four data sets. We finish with some discussions in Section 6.
@@ 2 Literature review
There have been numerous non-deep learning GRN inference methods, and they can be roughly classified as information-based and model-based. Information-based methods [12, 13, 8] do not rely on models for gene expression, but treat GRN inference as a feature selection problem: for a target gene, select genes that can be used to predict the level of the target gene. A common obstacle for information-based methods is to distinguish between direct and indirect regulations.
Model-based methods construct concrete models for gene expression under regulation, and fit the expression data to the models to determine model parameters and reconstruct the GRN. Some methods [14] require measuring the same cell multiple times, which is not quite applicable for now. Some methods [15, 16] use average expression levels measured over many cells (bulk level), which do not utilize the rich information in the single-cell level measurements. Some methods [17] only work on single-cell data at one time point. For single-cell gene expression data measured at multiple time points, where each cell is measured only once, we only know one model-based method, WENDY.
@@
Similar to TRENDY, there are some approaches to enhance GRNs inferred by other known methods[23, 24, 25, 26, 27], but they generally require extra knowledge of transcription factors or cannot fully determine the GRN.
Since WENDY paper and TRENDY paper contain more detailed summaries for traditional GRN inference methods and deep learning-based GRN inference methods, we present a review for GRN inference methods related to LLMs.
@@
Some researchers worked on pre-trained LLMs and fine-tuned them with new data. Weng et al. [31] trained GPT-3.5 with related papers to obtain GenePT, a new LLM that can provide a high-dimensional embedding (a vector of real numbers) of each gene. Then this embedding was used to train a neural network to output the GRN. Yang et al. [32] trained BERT with scRNAseq data to obtain scBERT, which can also provide gene embedding. Kommu et al. [33] used the embedding from scBERT to infer the GRN.
@@ 3.2 Algorithm details
In this section, given general covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ that may not satisfy Eqs. 13 for any $B_{0}$, we introduce a procedure to calculate $B$ that approximately solves Eqs. 13. Under some mild conditions (some matrices are invertible and have distinct eigenvalues), $B$ can be uniquely determined up to a $\pm$ sign. Then we use $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ to distinguish between $B$ and $-B$. The whole procedure is named as QWENDY method.
@@
With expression data from four time points, we uniquely determine $B$. See Algorithm 1 for the workflow of QWENDY method.
@@
Algorithm 1 Workflow of QWENDY method.
3.3 Correctness of QWENDY method
@@
4 Enhanced QWENDY method
QWENDY method is derived from Eqs. 16, especially Eqs. 13, which are a linear approximation of the actual nonlinear gene expression dynamics. Therefore, real $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ might not fit with Eqs. 13, and it might not be feasible to apply QWENDY directly to real $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$. This problem already exists for WENDY method, where the input matrices $K_{0}$, $K_{1}$ might not satisfy
@@ 4 Enhanced QWENDY method
For this problem, TRENDY method [10] proposes that we can construct $K_{1}^{*}=B^{\mathrm{T}}K_{0}B$, and train a model with input $K_{1}$ and target $K_{1}^{*}$, so that the output $K_{1}^{\prime}$ is close to $K_{1}^{*}$. Then we can apply WENDY to $K_{0}$, $K_{1}^{\prime}$ to obtain a more accurate GRN $A_{1}$. After that, we can train another model with input $K_{0}$, $K_{1}$, $A_{1}$, and the target is the true GRN $A_{\text{true}}$. Then the final output $A_{2}$ is more similar to the true GRN than $A_{1}$.
@@
$\mathrm{d}X_{j}(t)=V\left\{\beta\prod_{i=1}^{n}\left[1+(A_{\mathrm{true}})_{i,j}\frac{X_{i}(t)}{X_{i}(t)+1}\right]-\theta X_{j}(t)\right\}\mathrm{d}t+\sigma X_{j}(t)\mathrm{d}W_{j}(t),$ (12)
where $X_{i}(t)$ is the level of gene $i$ at time $t$, $W_{j}(t)$ is a standard Brownian motion, and $V=30$, $\beta=1$, $\theta=0.2$, $\sigma=0.1$ There are $10^{5}$ training samples, each with the expression levels of 10 genes for 100 cells, measured at four time points.
@@
5 Performance of different methods
The same as in the WENDY paper and the TRENDY paper [9, 10], we test different methods on two synthetic data sets and two experimental data sets. SINC data set [10] and DREAM4 data set [44] are generated by simulating stochastic differential equation systems. THP-1 data set is from monocytic THP-1 human myeloid leukemia cells [45]. hESC data set is from human embryonic stem cell-derived progenitor cells[46]. THP-1 and hESC data sets each has only one group of data (one ground truth GRN and the corresponding expression levels); DREAM4 data set has five groups of data; SINC data set has 1000 groups of data.
For SINC data set, we use data from all four time points. For DREAM4 data set, we use data from any four consecutive time points and take average. For THP-1 data set and hESC data set, we choose any four time points with equal difference. This is due to the limitation that QWENDY (and thus TEQWENDY and LEQWENDY) is derived for four evenly spaced time points.
We test the performance of QWENDY, TEQWENDY, and LEQWENDY methods on these four data sets. Since TRENDY ranks the first among 16 methods on these four data sets [10], we also present the performance of TRENDY. For each method, we compare the inferred GRN and the ground truth GRN by calculating the AUROC and AUPRC scores [9]. These two scores, both between 00 and $1$, evaluate the level of matching, where $1$ means perfect match, and 00 means perfect mismatch.
See Table 1 for the scores. We can see that TEQWENDY is slightly better than TRENDY, meaning that TEQWENDY has the best performance in all 19 methods (16 methods tested in TRENDY paper, QWENDY, TEQWENDY, and LEQWENDY). Besides, QWENDY is worse than TRENDY, and LEQWENDY is even worse than QWENDY, showing that LLMs trained with natural language inputs might not directly help with the numerical task in GRN inference.
Since TEQWENDY and TRENDY are trained on synthetic data, a natural concern is whether the training makes these methods overfitting to certain types of synthetic data, but less suitable for experimental data. Therefore, we also compare the performance of different methods only for two experimental data sets. For all 16 methods tested in TRENDY paper, bWENDY ranks the first on THP-1 and hESC data sets, with total score 1.5783 [10]. In comparison, QWENDY has total score 1.6272, making it the best in all 19 methods on two experimental data sets.
In sum, TEQWENDY method has the best overall performance, while QWENDY ranks the first on experimental data sets. LEQWENDY is worse than QWENDY in most cases, showing the failure of training with an LLM.
@@ 5 Performance of different methods
TRENDY
SINC AUROC 0.5314 0.4972 0.9587 0.8941
AUPRC 0.5720 0.5185 0.9142 0.8314
DREAM4 AUROC 0.4987 0.5164 0.5372 0.5341
AUPRC 0.1844 0.1823 0.2203 0.2177
THP-1 AUROC 0.5524 0.5543 0.5415 0.5557
AUPRC 0.4294 0.3632 0.3801 0.3669
hESC AUROC 0.6019 0.5905 0.4815 0.5311
AUPRC 0.0435 0.0367 0.0317 0.0376
Total 3.4137 3.2591 4.0652 3.9686
Table 1: AUROC and AUPRC scores of QWENDY, LEQWENDY, TEQWENDY, TRENDY methods on four data sets
6 Discussion
In this paper, we present QWENDY, a GRN inference method that requires single-cell gene expression data measured at four time points. Then we train the transformer-based variant TEQWENDY from scratch and fine-tune the LLM-based variant LEQWENDY based on pre-trained weights. TEQWENDY has the best overall performance, while QWENDY ranks the first on experimental data sets. Notice that each experimental data set only has one group of data, meaning that the performance of each method has a large uncertainty level.
The performance of LEQWENDY is not satisfactory, meaning that the pre-trained LLMs, even after task-specific fine-tuning, do not fit with this task. We need to explore other approaches of applying LLMs in GRN inference, and utilize the fact that LLMs are trained with natural language material.
QWENDY method and its variants require measurements at four evenly spaced time points $t_{0}$, $t_{1}$, $t_{2}$, $t_{3}$, meaning that $t_{1}-t_{0}=t_{2}-t_{1}=t_{3}-t_{2}$. Otherwise, the dynamics of covariance matrices is much more complicated, and we do not have an explicit solution. One possible strategy is to use the approximation
Since QWENDY and its variants need four time points, the duration of the whole experiment might be too long, so that the dynamics of gene expression under regulation might have changed during this time. QWENDY is based on a time-homogeneous model, which might fail in this situation.
Training TEQWENDY and LEQWENDY on data generated by Eq. 12 does not necessarily increase their performance on experimental data sets. This means that Eq. 12 might not faithfully reflect the gene expression dynamics. To better integrate deep learning techniques, we need better gene expression data generators.
Although QWENDY and its variants can fully determine the GRN, the diagonal elements that represent the regulation of one gene to itself (autoregulation) can be mixed with natural mRNA synthesis and degradation [47]. Therefore, it is not recommended to infer the autoregulation with QWENDY or its variants.
Given new GRN inference methods, we can study how cells maintain homeostasis [48, 49] or be driven away from homeostasis and cause diseases [50, 51].
Acknowledgments
Y.W. would like to thank Dr. Mingda Zhang for a helpful discussion.
Data and code availability
Main function of QWENDY, TEQWENDY, and LEQWENDY methods, including a tutorial and the training files, along with other data and code files used in this paper, can be found in
https://github.com/YueWangMathbio/QWENDY
Appendix: details of TEQWENDY and LEQWENDY

### v3

# QWENDY: Gene Regulatory Network Inference by Quadruple Covariance Matrices

@@ Abstract
Knowing gene regulatory networks (GRNs) is important for understanding various biological mechanisms. In this paper, we present a method, QWENDY, that uses single-cell gene expression data measured at four time points to infer GRNs. Based on a linear gene expression model, it solves the transformation of the covariance matrices. Unlike its predecessor WENDY, QWENDY avoids solving a non-convex optimization problem and produces a unique solution. We test the performance of QWENDY on two experimental data sets and two synthetic data sets. Compared to previously tested methods on the same data sets, QWENDY ranks the first on experimental data, although it does not perform well on synthetic data.
@@ 1 Introduction
One gene can activate or inhibit the expression of another gene. Genes and such regulation relations form a gene regulatory network (GRN). For $n$ genes, the corresponding GRN is commonly expressed as an $n\times n$ matrix $A$, where $A_{ij}>0/=0/<0$ means gene $i$ has positive/no/negative regulation effect on gene $j$, and its absolute value represents the regulation strength. If the regulation relations of concerned genes are known, one can understand the corresponding biological process or even control it with gene perturbation. Therefore, knowledge of GRNs can be useful in developmental biology [1, 2, 3], and even in the study of macroscopic behavior [4, 5, 6]. Since it is difficult to determine the GRN directly, the common practice is to infer the GRN from gene expression data. Given new GRN inference methods, we can study how cells maintain homeostasis [7, 8] or be driven away from homeostasis and cause diseases [9, 10].
@@
For such single-cell gene expression data measured at multiple time points, although it is the most informative data type under current technology, there are only a few GRN methods developed specifically for this data type [12]. We have developed the WENDY method to infer the GRN with this data type [13]. It needs measurement at two time points, and solves the transformation between the covariance matrices of two probability distributions, where the transformation is determined by the GRN. However, it needs to solve a non-convex optimization problem, which has infinitely many solutions, and WENDY will output one solution, determined by the numerical solver chosen in the realization.
In this paper, we present an improved version of WENDY that needs data measured at four time points, and it is also based on solving the transformation between covariance matrices. This new method is named QWENDY, where Q stands for quadruple. With the help of data from more time points, QWENDY can uniquely determine the GRN. We will prove that under some assumptions, the output of QWENDY is the ground truth GRN. Besides, QWENDY does not need to conduct non-convex optimizations, but just matrix decompositions. Therefore, the output of QWENDY does not depend on the realization of the numerical computing procedure.
In a previous paper [14], we have tested 16 GRN inference methods on two experimental data sets and two synthetic data sets. In this paper, we test the performance of QWENDY on the same data sets and compare it with previous methods. In all 17 GRN inference methods, QWENDY ranks the first on experimental data, although it does not perform well on synthetic data. These results suggest that QWENDY is a promising GRN inference method that warrants further testing and refinement.
Section 2 reviews other GRN inference methods. Section 3 introduces the QWENDY method, including a proof of its correctness. Section 4 tests the performance of QWENDY on four data sets. We finish with some discussions in Section 5.
@@ 2 Literature review
There have been numerous non-deep learning GRN inference methods, and they can be roughly classified as information-based and model-based.
Information-based methods [15, 16, 12] do not rely on models for gene expression, but treat GRN inference as a feature selection problem: for a target gene, select genes that can be used to predict the level of the target gene. A common obstacle for information-based methods is to distinguish between direct and indirect regulations.
Model-based methods construct concrete models for gene expression under regulation, and fit the expression data to the models to determine model parameters and reconstruct the GRN. Some methods [17] require measuring the same cell multiple times, which is not quite applicable for now. Some methods [18, 19] use average expression levels measured over many cells (bulk level), which do not utilize the rich information in the single-cell level measurements. Some methods [20] only work on single-cell data at one time point. For single-cell gene expression data measured at multiple time points, where each cell is measured only once, we only know one model-based method, WENDY. Therefore, we develop QWENDY as an alternative.
For more detailed summaries of traditional GRN inference methods, readers may refer to the literature review section of the WENDY paper [13].
@@
We have developed the TRENDY method [14] to enhance the WENDY method with deep learning tools. Similar to TRENDY, there are some approaches to enhance GRNs inferred by other known methods [26, 27, 28, 29, 30], but they generally require extra knowledge of transcription factors or cannot fully determine the GRN.
For more detailed summaries of deep learning-based GRN inference methods and approaches to enhance existing methods, readers may refer to the literature review section of the WENDY paper [13].
Recently, there are some GRN inference methods related to large language models (LLMs), although it is questionable whether LLMs trained with natural language data are useful to the study of gene regulation.
@@
Some researchers worked on pre-trained LLMs and fine-tuned them with new data. Weng et al. [34] trained GPT-3.5 with related papers to obtain GenePT, a new LLM that can provide a high-dimensional embedding (a vector of real numbers) of each gene. Then this embedding was used to train a neural network to output the GRN. Yang et al. [35] trained BERT with scRNA-seq data to obtain scBERT, which can also provide gene embedding. Kommu et al. [36] used the embedding from scBERT to infer the GRN.
@@ 3.2 Algorithm details
In this section, given general covariance matrices $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ that may not satisfy Eqs. 13 for any $B_{0}$, we introduce a procedure to calculate $B$ that approximately solves Eqs. 13. Under some mild conditions (some matrices are invertible and have distinct eigenvalues), $B$ can be uniquely determined up to a $\pm$ sign. Then we use $\mathbf{x}_{0}$, $\mathbf{x}_{1}$, $\mathbf{x}_{2}$, $\mathbf{x}_{3}$ to distinguish between $B$ and $-B$. The whole procedure is named as the QWENDY method.
@@
With expression data from four time points, we uniquely determine $B$. See Algorithm 1 for the workflow of the QWENDY method.
@@
Algorithm 1 Workflow of the QWENDY method.
3.3 Correctness of the QWENDY method
@@
4 Performance on synthetic and experimental data
The same as in the WENDY paper and the TRENDY paper [13, 14], we test the QWENDY method on two synthetic data sets and two experimental data sets. SINC data set [14] and DREAM4 data set [43] are generated by simulating stochastic differential equation systems. THP-1 data set is from monocytic THP-1 human myeloid leukemia cells [44]. hESC data set is from human embryonic stem cell-derived progenitor cells[45]. THP-1 and hESC data sets each has only one group of data (one ground truth GRN and the corresponding expression levels); DREAM4 data set has five groups of data; SINC data set has 1000 groups of data.
For SINC data set and DREAM4 data set, we use data from any four consecutive time points and take average. For THP-1 data set and hESC data set, we choose any four time points with equal difference. This is due to the limitation that QWENDY is derived for four evenly spaced time points.
We test the performance of the QWENDY method on these four data sets and compare it with the performance of 16 methods tested on the same data sets in the TRENDY paper [14]. We compare the inferred GRN and the ground truth GRN by calculating the AUROC and AUPRC scores [13]. These two scores, both between 00 and $1$, evaluate the level of matching, where $1$ means perfect match, and 00 means perfect mismatch.
See Table 1 for the performance of QWENDY, compared with other 16 previously tested methods. Although it does not perform very well on synthetic data sets, QWENDY has the best performance on experimental data sets.
QWENDY
Best in
other 16
Rank of QWE-
NDY in all 17
SINC AUROC 0.5107 0.8703 11th
AUPRC 0.5537 0.7672 7th
DREAM4 AUROC 0.4987 0.5741 11th
AUPRC 0.1844 0.2452 13th
Synthetic
total
1.7475 2.3584 11th
THP-1 AUROC 0.5524 0.6261 8th
AUPRC 0.4294 0.4205 1st
hESC AUROC 0.6019 0.6233 3rd
AUPRC 0.0435 0.0641 8th
Experimental
total
1.6272 1.5783 1st
Overall
total
3.3747 3.8497 9th
Table 1: AUROC and AUPRC scores of QWENDY on four data sets, compared with 16 previously tested methods
5 Discussion
In this paper, we present QWENDY, a GRN inference method that requires single-cell gene expression data measured at four time points. QWENDY ranks the first in 17 GRN inference methods on experimental data sets. Notice that each experimental data set only has one group of data, meaning that the performance of each method has a large uncertainty level.
The QWENDY method requires measurements at four evenly spaced time points $t_{0}$, $t_{1}$, $t_{2}$, $t_{3}$, meaning that $t_{1}-t_{0}=t_{2}-t_{1}=t_{3}-t_{2}$. Otherwise, the dynamics of covariance matrices is much more complicated, and we do not have an explicit solution. One possible strategy is to use the approximation
Since QWENDY needs four time points, the duration of the whole experiment might be too long, so that the dynamics of gene expression under regulation might have changed during this time. QWENDY is based on a time-homogeneous model, which might fail in this situation.
Although QWENDY can fully determine the GRN, the diagonal elements that represent the regulation of one gene to itself (autoregulation) can be mixed with natural mRNA synthesis and degradation [46]. Therefore, it is not recommended to infer the autoregulation with QWENDY or its variants.
Acknowledgments
Y.W. would like to thank Dr. Mingda Zhang for a helpful discussion.
Data and code availability
Data and code files used in this paper can be found in
https://github.com/YueWangMathbio/QWENDY
Appendix A Enhancing QWENDY with deep learning
A.1 Methods
WENDY method is derived from a simplified gene expression model, and the performance of WENDY is not satisfactory when the model does not fit with experiments. We developed the TRENDY method [14], which trains a transformer model, a deep learning architecture that can be applied to various fields [47], to transform the input data to better fit the gene expression model. Then the transformed data will produce more accurate GRNs by WENDY. Such more accurate GRNs will be further enhanced by another transformer model.
The QWENDY method is derived from Eqs. 16, especially Eqs. 13, which are a linear approximation of the actual nonlinear gene expression dynamics. Therefore, real $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$ might not fit with Eqs. 13, and it might not be feasible to apply QWENDY directly to real $K_{0}$, $K_{1}$, $K_{2}$, $K_{3}$. This problem already exists for the WENDY method, where the input matrices $K_{0}$, $K_{1}$ might not satisfy
@@ A.1 Methods
For this problem, the TRENDY method [14] proposes that we can construct $K_{1}^{*}=B^{\mathrm{T}}K_{0}B$, and train a model with input $K_{1}$ and target $K_{1}^{*}$, so that the output $K_{1}^{\prime}$ is close to $K_{1}^{*}$. Then we can apply WENDY to $K_{0}$, $K_{1}^{\prime}$ to obtain a more accurate GRN $A_{1}$. After that, we can train another model with input $K_{0}$, $K_{1}$, $A_{1}$, and the target is the true GRN $A_{\text{true}}$. Then the final output $A_{2}$ is more similar to the true GRN than $A_{1}$.
@@
$\mathrm{d}X_{j}(t)=V\left\{\beta\prod_{i=1}^{n}\left[1+(A_{\mathrm{true}})_{i,j}\frac{X_{i}(t)}{X_{i}(t)+1}\right]-\theta X_{j}(t)\right\}\mathrm{d}t+\sigma X_{j}(t)\mathrm{d}W_{j}(t),$ (S1)
where $X_{i}(t)$ is the level of gene $i$ at time $t$, $W_{j}(t)$ is a standard Brownian motion, and $V=30$, $\beta=1$, $\theta=0.2$, $\sigma=0.1$ There are $10^{5}$ training samples, each with the expression levels of 10 genes for 100 cells, measured at four time points: 0.0, 0.1, 0.2, 0.3.
@@
A.2 Performance
We also test the performance of TEQWENDY and LEQWENDY on the same data sets. See Table S1 for the scores. Compared with QWENDY, TEQWENDY performs better on synthetic data, but worse on experimental data. The overall performance of TEQWENDY is slightly better than QWENDY. Therefore, the improvement of TEQWENDY might be due to the overfitting on synthetic data, which does not work on experimental data. The performance of LEQWENDY is worse than QWENDY on both synthetic and experimental data, meaning that LLMs trained with natural language inputs might not directly help with the numerical task in GRN inference, even after task-specific fine-tuning.
Training TEQWENDY and LEQWENDY on data generated by Eq. S1 does not necessarily increase their performance on experimental data sets. This means that Eq. S1 might not faithfully reflect the gene expression dynamics. To better integrate deep learning techniques, we need better gene expression data generators. Another possibility is that the training data are only from time 0.0 - 0.3. We could use data from later time points to increase the robustness of training.
@@ A.2 Performance
SINC AUROC 0.5107 0.4990 0.5932
AUPRC 0.5537 0.5183 0.6014
DREAM4 AUROC 0.4987 0.5164 0.5372
AUPRC 0.1844 0.1823 0.2203
Synthetic
total
1.7475 1.7160 1.9521
THP-1 AUROC 0.5524 0.5543 0.5415
AUPRC 0.4294 0.3632 0.3801
hESC AUROC 0.6019 0.5905 0.4815
AUPRC 0.0435 0.0367 0.0317
Experimental
total
1.6272 1.5447 1.4348
Overall
total
3.3747 3.2607 3.3869
Overall rank
in all 19
10th 11th 9th
Table S1: AUROC and AUPRC scores of QWENDY, LEQWENDY, and TEQWENDY on four data sets. The overall rank is for 19 methods: 16 previously tested methods, QWENDY, LEQWENDY, and TEQWENDY
A.3 Technical details